At each point on the surface of the cube the electric field is parallel to the z axis. The length of each edge of the cube is 3.7 m. On the top face of the cube the electric field - 33 k N/C and on the bottom face it is + 22 k N/C. Determine the net charge contained within the cube.
Answer #1


flux through the top face phi 1= B1*A = 33*10^3*3.7^2 = 451770 Nm^2/C


flux through the bottom face phi 2 = B2*A = 22*10^3*3.7^2 = 301180 Nm^2/C


flux through front , back , left , right faces = 0

total flux = phi = 451770 + 301180 = 752950 Nm^2/C

from gauss law

total flux = Qin/eo

Qin = 752950*8.84*10^-12 = 6.7*10^-6 C